OpenStudy (anonymous):

Use the Pythagorean Theorem to find the missing side of the triangle. Round to the nearest tenth, if necessary. a = ?, b = 5, c = 8 A. 6.0 B. 6.1 C. 6.2 D. 6.3

OpenStudy (anonymous):

The Pythagorean Thm says that A^2 + B^2 = C^2. So we have: A^2 + 5^2 = 8^2 which is: A^2 + 25 = 64 A^2 = 39 A = sqrt(39)

OpenStudy (anonymous):